# Signals, Concepts, and Laws: Toward Universal, Explainable Time-Series Forecasting

### v2

@@ Abstract
Accurate, explainable and physically credible forecasting remains a persistent challenge for multivariate time-series with domain-varying statistical properties. We propose DORIC, a Domain-Universal, ODE-Regularized, Interpretable-Concept Transformer for Time-Series Forecasting that generates predictions through five self-supervised, domain-agnostic concepts while enforcing differentiable residuals grounded in first-principles constraints. Unlike prior efficiency-focused Transformers, such as Informer, which leverages sparsity, or FEDformer, which incorporates frequency priors, DORIC combines latent explainability with explicit scientific constraints, while preserving the attention mechanism’s capacity to model long-range dependencies. We evaluate DORIC on six publicly-available datasets and it achieves the lowest error in eight of twelve MSE/MAE metrics. Compared with TimeMixer, DORIC outperforms it on four datasets while maintaining strong explainability. Ablation studies reveal complementary contributions: removing the physics residual increases average MSE from 0.328 to 0.547, eliminating concept alignment raises it to 0.698, and replacing the shared encoder with disjoint concept heads results in a 76% increase.
@@ 1 Introduction
Real-world decision systems, from power-grid dispatch and urban mobility control to epidemic surveillance and high-frequency trading, depend on fast, accurate, and trustworthy forecasts of multivariate time-series. Classical linear models, exemplified by the Box-Jenkins ARIMA methodology (Tunnicliffe Wilson, 2016), provide statistical rigor but struggle with high-frequency, non-stationary data or dynamics governed by latent physical constraints. Over the past decade, attention-based deep learning has revolutionized sequential modelling: the Transformer architecture introduced by Vaswani et al. (2017) replaced recurrence with self-attention, enabling parallel learning over long-contexts. Subsequent efficiency-oriented variants, such as Reformer (Kitaev et al., 2020) and LogTrans (Li et al., 2019), reduce memory overhead via locality-sensitive hashing and log-sparse kernels. Meanwhile, domain-specific advances, such as Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), and FEDformer (Zhou et al., 2022), introduced decomposition, auto-correlation and frequency sparsity to enhance long-horizon accuracy. More recently, PatchTST (Nie et al., 2023b) demonstrated further gains via patching and exponential-smoothing priors. A comprehensive survey (Wen et al., 2023) attests to the field’s explosive growth. Despite these advancements, three structural gaps remain.
@@
  • Gap 2 – Latent opacity. Although post-hoc explanation methods exist, their fidelity degrades in multivariate settings. In contrast, concept-driven transparency pioneered by Concept Bottleneck Models (Koh et al., 2020) and their stochastic extensions (Vandenhirtz et al., 2024) remains under-explored in the context of time-series forecasting.
@@
To bridge these gaps, we propose DORIC (Regularized Conceptual Neural Networks with Attention) — an end-to-end framework that: (i) routes all information through five novel, domain-agnostic concepts capturing level, growth, periodicity, volatility and exogenous pressure; (ii) embeds a physics-informed residual loss derived from first-principles conservation equations; and (iii) retains the quadratic-free ProbSparse attention block to ensure scalability on very long sequences. Extensive experiments on six public benchmarks, Electricity, Traffic, Weather, Illness, Exchange-rate and ETT, show that DORIC consistently outperforms competitive baselines, including FF Bottleneck, Reformer, LogTrans, Informer, Autoformer, FEDformer and TimeMixer (van Sprang et al., 2025; Kitaev et al., 2020; Li et al., 2019; Zhou et al., 2021; Wu et al., 2021; Zhou et al., 2022; Wang et al., 2024), and a strong classical AR model, achieving substantial reductions in both mean squared error (MSE) and mean absolute error (MAE).
The major contributions of our research can be summarized as follows:
  1. 1.
    Concept-centred attention. We design a learnable bottleneck that projects raw multivariate inputs onto five interpretable, scale-free concepts without requiring external labels. This design enables instantaneous attribution and interpretability, eliminating the need for post-hoc processing.
  2. 2.
    Physics-regularised loss. A residual term penalises deviations from domain-specific conservation laws, guiding optimization toward physically plausible states while maintaining differentiability for end-to-end backpropagation.
  3. 3.
    Domain-universal architecture. Leveraging adaptive positional encoding and shared concept heads, DORIC is trained once and transfers effectively across diverse domains, including electricity load curves, traffic occupancy, meteorological records, epidemiological counts, currency pairs and industrial benchmarks with only minimal hyper-parameter adjustments.
@@ 2 Literature Review
Early time‑series forecasting relies on stochastic linear models such as the Box–Jenkins ARIMA (Tunnicliffe Wilson, 2016), whose identification–estimation–diagnosis cycle remains influential but often fails under the non‑stationarity of modern telemetry.
@@
Deep learning models capture time-series patterns with specifically-designed architectures spanning a wide range of foundation backbones, including CNNs (Wang et al., 2023; Wu et al., 2023; Hewage et al., 2020), RNNs (Lai et al., 2018; Qin et al., 2017; Salinas et al., 2020), Transformers (Vaswani et al., 2017) and MLPs (Zeng et al., 2023b; Zhang et al., 2022; Oreshkin et al., 2019; Challu et al., 2023). The quadratic cost of vanilla attention spurred efficient variants, Reformer (Kitaev et al., 2020) and LogTrans (Li et al., 2019), and domain‑specific advances: Informer’s ProbSparse attention (Zhou et al., 2021), Autoformer’s trend–seasonality decomposition (Wu et al., 2021), FEDformer’s Fourier sparsity (Zhou et al., 2022), PatchTST’s patch‑token strategy (Nie et al., 2023b), and TimeMixer (Wang et al., 2024). A 2023 survey (Wen et al., 2023) documents more than 120 Transformer adaptations for time‑series forecasting.
@@
Despite these advances, interpretability remains a challenge, prompting growing interest in concept-level supervision. Concept Bottleneck Models (Koh et al., 2020) demonstrate that forcing predictions through human-meaningful variables can maintain accuracy while enabling user interventions. Stochastic CBMs (Vandenhirtz et al., 2024) generalized this idea by modelling concept dependencies probabilistically. FF Bottleneck (van Sprang et al., 2025) used an AR model as a surrogate concept within a Transformer architecture; however, empirical results revealed that the standalone AR model outperformed FF Bottleneck on four out of the six datasets.
@@
In parallel, physics-informed learning has emerged as a remedy for physically implausible outputs. Raissi’s PINNs framework (Raissi et al., 2019) pioneered the joint optimization of data and governing equations; PGTransNet (Wu et al., 2024) integrated physics-guided self-attention for Pacific-Ocean temperature forecasting, while physics-informed LSTM variants have been applied to power-transformer health monitoring (Lei et al., 2025). The recently proposed PhysicsSolver (Zhu et al., 2025) further attests to the potential of incorporating scientific priors into deep learning architectures.
@@
Hybrid work now marries physics with deep networks: a TCN–TFT ensemble improves probabilistic wind forecasts (Mi et al., 2025), while energy‑constrained diffusion Transformers extend long‑horizon accuracy (Ren et al., 2024). Interpretability has advanced through Temporal Fusion Transformer’s gating and variable‑selection (Lim et al., 2021) and through sparse‑attention schemes such as Adversarial Sparse Transformer (Wu et al., 2020) and Query‑Selector attention (Klimek et al., 2022). Complementing advances in model architecture, evaluation paradigms have matured: GraphCast — a graph-neural extension that outperforms ECMWF simulations on 90% of medium-range weather targets (Lam et al., 2023) —demonstrates that ML systems can rival numerical solvers at global scale. Meanwhile, foundation models like TimeGPT (Garza et al., 2024) deliver strong zero-shot generalization capabilities across diverse time-series domains. Nevertheless, critical evaluations have shown that simple linear baselines can outperform many Transformer variants when evaluated rigorously, as highlighted in the “Are Transformers Effective for Time-Series Forecasting?” paper (Zeng et al., 2023a). Additionally, the rise of diffusion-based Transformers (Fu et al., 2025) reflects the field’s continued pursuit of the right balance between model complexity and effective inductive biases.
@@ 3.2.1 Initial Token Embedding:
where $\mathbf{W}_{e}\in\mathbb{R}^{1\times d}$ (row) is the embedding matrix, mapping each scalar value to a $d$ dimensional vector (in experiments, we choose $d=64$), $\mathbf{P}\in\mathbb{R}^{L\times d}$ encodes absolute positions via sinusoidal basis: $P_{i,2k}=\sin(i/10^{4k/d})$, $P_{i,2k+1}=\cos(i/10^{4k/d})$, and $\mathbf{H}^{(0)}\in\mathbb{R}^{L\times d}$ is the initial token matrix.
@@ 3.3.2 Analytical Soft Targets:
We endow those five coordinates with physical meaning through soft targets
@@
generated by causal statistics:
@@ 3.3.3 Concept Alignment Loss:
$\mathcal{L}_{\text{concept}}=\frac{1}{N}\sum_{t=L+1}^{T}\bigl\|\mathbf{c}_{t}-\mathbf{c}_{t}^{\star}\bigr\|_{2}^{2}.$
Why soft, not hard?
Soft supervision allows the model robust to missing values, noise, cross-column differences, and can adaptively adjust the optimal window length.
@@ 3.5 Joint Loss
We fix $(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4},\lambda_{y})=(1,1,1,1,1)$, physics penalty $\lambda_{\text{phys}}$ and the concept penalty $\lambda_{\text{con}}$ are both 1.
@@ 4 Experiments and Analysis
4.1 Datasets and Codes
To establish the practical value of DORIC we conduct a comprehensive evaluation on six widely-used public benchmarks that span very different sampling frequencies, signal-to-noise ratios, and seasonality regimes. For more information on the datasets, we refer the reader to Appendix A.
@@ 4.2 Baselines and Implementation Details of DORIC
(1) AR – classic autoregressive linear model (order selected by AIC). (2) FF Bottleneck (van Sprang et al., 2025) – Transformer with a surrogate AR concept layer. (3) LogTrans (Li et al., 2019), Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), FEDformer (Zhou et al., 2022), PatchTST (Nie et al., 2023a), TimeMixer (Wang et al., 2024), – state-of-the-art Transformer variants representative of locality-sensitive hashing, log-sparsity, ProbSparse, decomposition autocorrelation, and Fourier sparsity, respectively.
@@ 4.3 Quantitative Results and Analysis
Quantitative performance :
Traffic: DORIC achieves MSE 0.313 and MAE 0.226, clearly better than TimeMixer/PatchTST (both at 0.360 MSE). In this high-variance regime, the growth/volatility concepts + physics residuals help suppress bursty errors and reduce overall risk.
Exchange rate: MSE 0.051, MAE 0.168—best among deep models. The combination of local fluctuation + periodicity + a forced-damped ODE prior aligns well with multi-scale currency dynamics.
Illness: MSE 0.869, MAE 0.740—best. DORIC captures structural breaks induced by public-health interventions; a mild positive bias suggests enriching the “exogenous pressure” concept with calendar/policy signals.
ETT: MSE 0.111, MAE 0.236, superior to most deep baselines (though the linear AR baseline is unusually strong here). This dataset couples quasi-physical device behavior with strong periodicity, indicating benefits from adaptive physical-loss scheduling.
Weather: DORIC is nearly tied with AR (MSE 0.007 vs. 0.006) and clearly better than deep baselines. Over-weighting the physical term on humidity channels can cause amplitude shrinkage, implying a need for softer/learned physical weights in “soft-law” variables.
Electricity: MAE is best (0.214) while MSE is slightly higher (0.138 vs. 0.129 for PatchTST/TimeMixer). This pattern indicates smaller routine errors with conservative handling of extreme spikes—consistent with the forced-damped ODE and variance-kinematics residuals that smooth outliers (good for MAE, sometimes raising MSE on a few peaks).
Summary: DORIC’s cross-domain stability stands out. Aside from specific linear channels where AR can be a near-upper-bound, DORIC consistently beats deep baselines while preserving interpretability and physical plausibility.
Why does it work?
Five-concept bottleneck—level, growth, power, dominant-period amplitude, local volatility—makes representations alignable and softly supervised, mitigating the opacity of pure attention and allowing robust correction under noise/missingness.
Physics residuals encode concept–observable relations as differentiable constraints (e.g., mean-derivative matches “velocity”; forced-damped ODE on the target), restricting the hypothesis space to a physically feasible manifold. Training tends to fit physics first, then refine with data and concepts.
4.4 Fitting and Prediction Analysis

### v3

@@ Abstract
Accurate, explainable and physically credible forecasting remains a persistent challenge for multivariate time-series with domain-varying statistical properties. We propose DORIC, a Domain-Universal, ODE-Regularized, Interpretable-Concept Transformer for Time-Series Forecasting that generates predictions through five self-supervised, domain-agnostic concepts while enforcing differentiable residuals grounded in first-principles constraints. The concepts are softly regressed toward analytic statistics of the raw signal, and a driven–damped ODE head couples these concepts to the forecast as a shared, mean-reverting dynamical template across datasets. Unlike prior efficiency-focused Transformers, such as Informer(sparse attention) or FEDformer(frequency priors), DORIC combines latent explainability with explicit scientific constraints, while preserving the attention mechanism’s capacity to model long-range dependencies. We evaluate DORIC on six publicly-available datasets and it achieves the lowest error in eight of twelve MSE/MAE metrics. Compared with TimeMixer, DORIC outperforms it on four datasets while maintaining strong interpretability. Interpretability analyses show that the learned concepts remain strongly aligned with their analytic targets, physics residuals stay relative to the signal scale, and the learned ODE coefficients follow domain-consistent patterns. Ablation studies reveal complementary contributions: removing the physics residual increases average MSE from 0.328 to 0.547, eliminating concept alignment raises it to 0.698, and replacing the shared encoder with disjoint concept heads results in a 76% increase.
@@ 1 Introduction
Real-world decision systems, from power-grid dispatch and urban mobility control to epidemic surveillance and high-frequency trading, depend on fast, accurate, and trustworthy forecasts of multivariate time-series. Classical linear models, exemplified by the Box-Jenkins ARIMA methodology (2), provide statistical rigor but struggle with high-frequency, non-stationary data or dynamics governed by latent physical constraints. Over the past decade, attention-based deep learning has revolutionized sequential modelling: the Transformer architecture introduced by 1 replaced recurrence with self-attention, enabling parallel learning over long-contexts. Subsequent efficiency-oriented variants, such as Reformer (3) and LogTrans (4), reduce memory overhead via locality-sensitive hashing and log-sparse kernels. Meanwhile, domain-specific advances, such as Informer (5), Autoformer (6), and FEDformer (7), introduced decomposition, auto-correlation and frequency sparsity to enhance long-horizon accuracy. More recently, PatchTST (8) demonstrated further gains via patching and exponential-smoothing priors. A comprehensive survey (10) attests to the field’s explosive growth. Despite these advancements, three structural gaps remain.
@@
  • Gap 2 – Latent opacity. Although post-hoc explanation methods exist, their fidelity degrades in multivariate settings. In contrast, concept-driven transparency pioneered by Concept Bottleneck Models (11) and their stochastic extensions (12) remains under-explored in the context of time-series forecasting.
@@
To bridge these gaps, we view long-horizon forecasting through an interpretability-first lens. Rather than adding ever more opaque modules, we ask whether a single concept–physics layer can serve as a reusable interface between diverse multivariate time series and human experts. Concretely, we introduce DORIC, a forecasting architecture that augments a standard encoder with (i) a small set of physically motivated concepts computed from the input signals, and (ii) a physics-informed head that couples these concepts via simple laws. This yields a model whose internal “dials” have stable semantics across datasets while remaining competitive with strong baselines.
Contributions.
The contributions of this work are threefold:
  • A universal concept–physics layer for multivariate time series. We design five low-level yet broadly applicable concepts (level, velocity, instantaneous power, dominant periodic amplitude, and local volatility) that can be attached to standard sequence encoders. These concepts are computed by causal statistics and serve as a shared vocabulary across datasets, rather than being tailored to any single benchmark.
  • A physics-informed head with provable training dynamics. We couple the concepts to the forecast through a driven–damped first-order ODE and turn this ODE into algebraic residuals. The resulting physics loss admits a ramp-up schedule under which SGD converges while the physics violation vanishes. This connects the training trajectory to a clear three-stage story: feasibility (physics residual collapse), concept alignment, and data fit.
  • An interpretability-first evaluation of forecasting performance. We show on seven standard benchmarks that plugging the concept–physics layer into a strong Transformer backbone achieves accuracy comparable to recent state-of-the-art models, while enabling concept-level analyses, gradient-based local sensitivities, and cross-dataset comparisons. Ablations demonstrate how concept supervision and physics penalties jointly control the trade-off between predictive accuracy and interpretability.
Scope of interpretability and scientific value.
Our interpretability claim is scoped at the level of the shared bottleneck and dynamical template rather than at every individual weight in the network. Each forecast $\hat{y}_{t}$ is required to pass through a five-dimensional concept vector $c_{t}=(c_{1,t},\dots,c_{5,t})$ and a driven–damped ODE head that are shared across all datasets. These coordinates are not arbitrary hidden units: they are softly regressed towards analytic statistics of the raw signal (sliding mean, local velocity, instantaneous power, dominant periodic amplitude, local volatility), so that the bottleneck remains tied to physically meaningful quantities instead of drifting into opaque features. The ODE head then combines these concepts through a mean-reverting dynamics that is also shared across domains. Together, this design yields a mediating mechanism from past observations to future predictions: time points influence the forecast only through how they shape the concept trajectories and how those trajectories are processed by the ODE. We do not claim that these five concepts exhaust all possible explanations, but that DORIC enforces a low-dimensional, physically anchored interface that is consistent across heterogeneous time-series domains.
@@ 2 Literature Review
Early time‑series forecasting relies on stochastic linear models such as the Box–Jenkins ARIMA (2), whose identification–estimation–diagnosis cycle remains influential but often fails under the non‑stationarity of modern telemetry.
@@
Deep learning models capture time-series patterns with specifically-designed architectures spanning a wide range of foundation backbones, including CNNs (32; 33; 34), RNNs (35; 36; 37), Transformers (1) and MLPs (38; 39; 40; 41). The quadratic cost of vanilla attention spurred efficient variants, Reformer (3) and LogTrans (4), and domain‑specific advances: Informer’s ProbSparse attention (5), Autoformer’s trend–seasonality decomposition (6), FEDformer’s Fourier sparsity (7), PatchTST’s patch‑token strategy (8), and TimeMixer (30). A 2023 survey (10) documents more than 120 Transformer adaptations for time‑series forecasting.
@@
Despite these advances, interpretability remains a challenge, prompting growing interest in concept-level supervision. Concept Bottleneck Models (11) demonstrate that forcing predictions through human-meaningful variables can maintain accuracy while enabling user interventions. Stochastic CBMs (12) generalized this idea by modelling concept dependencies probabilistically. FF Bottleneck (9) used an AR model as a surrogate concept within a Transformer architecture; however, empirical results revealed that the standalone AR model outperformed FF Bottleneck on four out of the six datasets.
@@
In parallel, physics-informed learning has emerged as a remedy for physically implausible outputs. Raissi’s PINNs framework (13) pioneered the joint optimization of data and governing equations; PGTransNet (15) integrated physics-guided self-attention for Pacific-Ocean temperature forecasting, while physics-informed LSTM variants have been applied to power-transformer health monitoring (16). The PhysicsSolver (17) further attests to the potential of incorporating scientific priors into deep learning architectures.
@@
Hybrid work now marries physics with deep networks: a TCN–TFT ensemble improves probabilistic wind forecasts (18), while energy‑constrained diffusion Transformers extend long‑horizon accuracy (19). Interpretability has advanced through Temporal Fusion Transformer’s gating and variable‑selection (20) and through sparse‑attention schemes such as Adversarial Sparse Transformer (21) and Query‑Selector attention (22). Complementing advances in model architecture, evaluation paradigms have matured: GraphCast — a graph-neural extension that outperforms ECMWF simulations on 90% of medium-range weather targets (23) —demonstrates that ML systems can rival numerical solvers at global scale. Meanwhile, foundation models like TimeGPT (24) deliver strong zero-shot generalization capabilities across diverse time-series domains. Nevertheless, critical evaluations have shown that simple linear baselines can outperform many Transformer variants when evaluated rigorously, as highlighted in the “Are Transformers Effective for Time-Series Forecasting?” paper (25). Additionally, the rise of diffusion-based Transformers (26) reflects the field’s continued pursuit of the right balance between model complexity and effective inductive biases.
@@ 3.2.1 Initial Token Embedding:
where $\mathbf{W}_{e}\in\mathbb{R}^{1\times d}$ (row) is the embedding matrix, mapping each scalar value to a $d$ dimensional vector, $\mathbf{P}\in\mathbb{R}^{L\times d}$ encodes absolute positions via sinusoidal basis: $P_{i,2k}=\sin(i/10^{4k/d})$, $P_{i,2k+1}=\cos(i/10^{4k/d})$, and $\mathbf{H}^{(0)}\in\mathbb{R}^{L\times d}$ is the initial token matrix.
@@ 3.3.2 Analytical Soft Targets:
We endow those five coordinates with physical meaning through soft targets generated by causal statistics:
@@
We denote by $y_{t}$ the target series at time $t$ and by $X_{t}\in\mathbb{R}^{N}$ the full multivariate input at time $t$. From each history segment $y_{t-\tau:t-1}$ we derive a vector of concept targets $c_{t}^{\star}=(c_{1,t}^{\star},\dots,c_{5,t}^{\star})\in\mathbb{R}^{5}$ using simple causal statistics. These $c_{t}^{\star}$ serve as soft targets for the concept predictor $g_{\phi}\circ f_{\theta}$; the network is not forced to exactly reproduce every statistic, but is nudged toward interpretable regions of the feature space.
@@
Rather than assigning a separate semantic label to every time index, DORIC explains forecasts through the evolution of a small set of shared concepts and their contributions in the ODE head. Individual time points influence the prediction only via how they update the concepts, which we argue is a more stable and transferable interface than attempting to label every $y_{t}$ directly.
Notation for concepts and soft targets.
At each time $t$ we have a five-dimensional concept vector $c_{t}=(c_{1,t},\dots,c_{5,t})\in\mathbb{R}^{5}$ and an analytic “soft target” $c_{t}^{\ast}=(c_{1,t}^{\ast},\dots,c_{5,t}^{\ast})\in\mathbb{R}^{5}$ defined by Eqs. (10)–(11). We use the following convention:
  • $c_{1,t}$ (level): sliding mean of $y$ over a causal window of length $\tau$;
  • $c_{2,t}$ (growth): local velocity (finite difference of the level);
  • $c_{3,t}$ (instantaneous power): local energy, proportional to $y_{t}^{2}$;
  • $c_{4,t}$ (dominant periodic amplitude): magnitude of the leading seasonal harmonic in the window;
  • $c_{5,t}$ (local volatility): standard deviation of the detrended signal over the same window.
@@ 3.3.3 Concept Alignment Loss:
The concept alignment loss then penalizes deviations between the predicted concepts $c_{t}=g_{\phi}(f_{\theta}(X_{t-L+1:t}))$ and the causal statistics $c_{t}^{\star}$:
$\mathcal{L}_{\text{concept}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=L+1}^{T}\big\|c_{i,t}-c_{i,t}^{\star}\big\|_{2}^{2}.$ (12)
We treat $c_{t}^{\star}$ as soft supervision rather than hard constraints. This makes the model robust to missing values, noise, and cross-column heterogeneity, and lets the network adjust the effective window length and weighting within each statistic when this yields better data fit without destroying the interpretability of each concept.
@@ 3.4.1 Driven-damped ODE:
Role of the ODE template.
Importantly, we do not assume that the data-generating process is exactly governed by the linear driven–damped ODE. Instead, the ODE acts as a shared, high-level dynamical template that biases the one-step forecaster $h_{\psi}$ towards mean-reverting, concept-driven behaviour. The residual $R_{y}(t)$ measures the inconsistency between the prediction and this template and enters the loss as a soft penalty, rather than a hard constraint. This design allows DORIC to benefit from a physically motivated regularizer while retaining enough flexibility to fit complex real-world dynamics.
@@ 4 Experiments and Analysis
4.1 Datasets and Further Analyses
To establish the practical value of DORIC we conduct a comprehensive evaluation on six widely-used public benchmarks that span very different sampling frequencies, signal-to-noise ratios, and seasonality regimes. For more information on the datasets and further analyses, we refer the reader to Appendix.
@@ 4.2 Baselines and Implementation Details of DORIC
(1) AR – classic autoregressive linear model (order selected by AIC). (2) FF Bottleneck (9) – Transformer with a surrogate AR concept layer. (3) LogTrans (4), Informer (5), Autoformer (6), FEDformer (7), PatchTST (29), TimeMixer (30), – state-of-the-art Transformer variants representative of locality-sensitive hashing, log-sparsity, ProbSparse, decomposition autocorrelation, and Fourier sparsity, respectively.
@@ 4.3 Quantitative Results
4.4 Concept–increment correlations
Figure 2 visualizes Pearson correlations between $\Delta y_{t}$ and each concept $c_{k,t}$ for representative channels. Two robust trends emerge:
Sanity checks for interpretability. We recommend three inexpensive audits per dataset: (i) local sensitivity $\partial\hat{y}/\partial c_{k}$ at typical points (should match the sign logic of the ODE head); (ii) counterfactual nudges $c_{k}\!\mapsto\!c_{k}+\delta$ (small $\delta$) with other concepts frozen, verifying that trajectories evolve consistently with residual identities; (iii) time-consistency of concept statistics (e.g., $c_{1}$ tracks sliding mean; $c_{5}$ tracks local volatility).
Figure 2: Concept Correlation Heatmap.
4.5 Prediction and Interpretability Analysis